How do you find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle?
Answer by stanbon(26291)

How do you find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle?
-------------------
Let the angle be "x".
The complement is "90-x"
The supplement is "180-x"
------------------------------------
Equation:
3(90-x) = 180-x + 50
270 - 3x = 230 -x
2x = 40
x = 20 degrees (the angle)
=============================
Cheers,
Stan H.